[Hallc_running] [New Logentry] RC Report -- 09/04/22

dhamilto at jlab.org dhamilto at jlab.org
Sun Sep 4 10:10:01 EDT 2022

Logentry Text:
We are currently in production running at 2.5 uA on LD2 for the final kinematic setting (Q2=1.60, W=3.08 Th_shms=6.00). We have 2347 pi- events and have collected 290 mC, which is approaching half of the required amount of data. Aside from a couple of short accelerator down periods yesterday as Hall B was coming up, there were no issues (18.4 ABU over the last three shifts).

If conditions continue as they have been, it is likely that this setting will be finished before Friday. If so, we are thinking of returning to the Q2 = 3.85 GeV2 setting for which the stats are somewhat low. There is also the low current BCM calibration making use of the Faraday cup to be scheduled, perhaps immediately after the maintenance day on Wednesday.

Since there is not much to discuss I propose we skip the RC meeting again today. We should probably meet tomorrow to discuss plans for later in the week.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4037704
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220904/03685c87/attachment.html>

More information about the Hallc_running mailing list